A medical researcher wants to investigate the amount of times it takes for patients’ headache pain to be relived after taking a new prescription painkiller. She plans to use statistical methods to estimate the mean of the population of relief times. She believes that the population is normally distributed with a standard deviation of 23 minutes. How large a sample should she take to estimate the mean time to within 3 minutes with 99% confidence?
Answer by Boreal(15235)

half-interval is 3 minutes=z(0.995)*sigma/sqrt(n), z=2.576 sigma is 23
so 3=2.576*23/sqrt(n)
sqrt(n)=2.5758*23/3=19.749
n=390 people
note, if one uses 2.576 for z, the answer it 390.03 and that would be rounded up to 391.